# Cascading from the winds to the disk: the universality of supernovae-driven turbulence in different galactic interstellar mediums

### v1

@@ Acknowledgments
We thank Norm Murray, Peng Oh, Jim Stone, Ralf Klessen, Christoph Federrath, the CITA plasma-astro group for the many enlightening discussions. I. C. and J. R. B. acknowledge compute allocations rrg-ripperda and rrg-essick from the Digital Research Alliance of Canada, which were utilized for both running and analyzing the simulations. I. C. acknowledges support from the Barry M. Goldwater Scholarship and acknowledges funding from the UC Santa Cruz Undergraduate Research in Science & Technology (URST) award. J. R. B. acknowledges funding from the Natural Sciences and Engineering Research Council of Canada (NSERC) (funding reference number 568580) and acknowledges the support from NSF Award 2206756, as well as high-performance computing resources provided by the Leibniz Rechenzentrum and the Gauss Center for Supercomputing grants pn76gi pr73fi and pn76ga. This research was made possible thanks to funding from the Lamat Institute and UC Santa Cruz through the Heising-Simons Foundation and NSF grants: AST-1852393, AST-1911206, AST-2150255 and AST-2206243.
@@ Appendix A Error propagation and fitting the galactic profile models
[…]
@@
and $a_{1}$ is the acceleration of the stellar disk. Solving,
@@
Appendix B Divergence and Compressible Mode Power Spectrum
In Section 4 we show the compressible mode $\bm{u}_{c}$ spectrum was directly related to the sound speed spectrum, which was controlling the ISM phase structure in our models with a thick disk (Milky Way analogue models). Here we formalize the connection with $\bm{u}_{c}$ and the fluid divergence. The divergence $\bm{\nabla}\cdot\bm{u}$ and $\bm{u}_{c}$ are intrinsically related because only $\bm{u}_{c}$ satisfies $\bm{\nabla}\cdot\bm{u}\neq 0$. In Figure 10 we plot the power spectrum of both quantities (left MW models, right SB models), with the expectation that $u_{c}^{2}(k)\propto(\bm{\nabla}\cdot\bm{u})^{2}(k)/k^{2}$ based on dimensionality. Indeed, we find that $u_{c}^{2}(k)\propto(\bm{\nabla}\cdot\bm{u})^{2}(k)/k^{2}$ across all resolved modes in the simulations. As we show in Figure 7, $u_{c}^{2}(k)\propto k^{-2}$, which we associate with Burgers (1948) turbulence (as in Beattie et al. 2025a on large scales).
@@ Appendix B Divergence and Compressible Mode Power Spectrum
Figure 11: One-dimensional, isotropic velocity power spectra, $u^{2}(k)$, normalized by the integrated power, as a function of wavenumber, $k$, normalized by the system scale wavenumber, $2\pi/L$. We compensate the spectra by $k^{-5/3}$ to compare to the Kolmogorov 1941 power spectrum. MW runs are shown on the left panel and SB on the right, with each color indicating a different grid resolution of the FX (pinks) and SC (purples) seeding schemes.
Appendix C Convergence test

### v2

@@ Acknowledgments
We thank the anonymous reviewer for helping increase the clarity and presentation of the study. We further thank Norm Murray, Peng Oh, Jim Stone, Ralf Klessen, Christoph Federrath, and the CITA plasma-astro group for the many enlightening discussions. I. C. and J. R. B. acknowledge compute allocations rrg-ripperda and rrg-essick from the Digital Research Alliance of Canada, which were utilized for both running and analyzing the simulations. I. C. acknowledges support from the Barry M. Goldwater Scholarship and acknowledges funding from the UC Santa Cruz Undergraduate Research in Science & Technology (URST) award. J. R. B. acknowledges funding from the Natural Sciences and Engineering Research Council of Canada (NSERC) (funding reference number 568580) and acknowledges the support from NSF Award 2206756, as well as high-performance computing resources provided by the Leibniz Rechenzentrum and the Gauss Center for Supercomputing grants pn76gi, pr73fi and pn76ga. This research was made possible thanks to funding from the Lamat Institute and UC Santa Cruz through the Heising-Simons Foundation and NSF grants: AST-1852393, AST-1911206, AST-2150255 and AST-2206243.
@@ Appendix A Error propagation and fitting the galactic profile models
We model the vertical density profile of the disk using Equation 17 (detailed in Section 3.2) and fit the free parameter $\sigma_{z}^{2}$ (Equation 16). We employ a MCMC maximum-likelihood method using the emcee (Foreman-Mackey et al., 2013) ensemble sampler to construct the posterior. The likelihood was constructed from the log-density profiles $\log\rho(z)$, weighted by the variance at each height $z$, and an additional Gaussian window function to extract only the disk component. We pick uniform priors: $10^{4}\,\rm{cms}^{-1}<\sigma_{z}<8\times 10^{6}\,\rm{cms}^{-1}$, based on physical constraints for the turbulent and thermally supported disk. For each model we ran $50$ chains for 5,0005,\!000 steps, discarding the first $20\%$ of the chains as burn-in stage. We report the median and standard deviation of $\sigma_{z}$ as the best-fit value and uncertainty directly from the posterior. From $\sigma_{z}$ and Equation 18, we derive the gaseous scale height $\ell_{0}$, and propagate the uncertainty by applying standard Gaussian error propagation, where
@@
and $a_{1}$ is the acceleration of the stellar disk. Putting this together,
@@
Appendix B Fitted Power Law Indexes For Spectra
Table 3: Fitted power laws of the one-dimensional, isotropic power spectra quantities normalized by total power, as a function of wavenumber, $k$, normalized by the system scale wavenumber, $2\pi/L$, in Figure 7 for each model.
Model $du^{2}(k)/dk$ $du_{s}^{2}(k)/dk$ $du_{c}^{2}(k)/dk$ $d\rho u^{2}(k)/dk$ $dc_{s}^{2}(k)/dk$
(1) (2) (3) (4) (5) (6)
MWSC $1.48\pm 0.12$ $1.30\pm 0.11$ $94\pm 0.17 1.1$$16\pm 0.05$ $1.83\pm 0.10$
MWFX $1.43\pm 0.14$ $1.26\pm 0.12$ $95\pm 0.21 1.1$$19\pm 0.04$ $1.73\pm 0.12$
SBSC $1.24\pm 0.05$ $1.07\pm 0.05$ $1.73\pm 0.12$ $0.39\pm 0.14$ $1.20\pm 0.07$
SBFX $1.14\pm 0.04$ $1.01\pm 0.03$ $1.59\pm 0.10$ $0.61\pm 0.03$ $1.19\pm 0.03$
Note. — Column (1): Numerical model label, where MW means Milky Way analogue, SB means starburst analogue, SC refers to the self-consistent SNe seeding prescription, and FX refers to the fixed seeding prescription (see Section 2.2.2 for more details about the exact seeding prescription). All indexes are derived from fits across the range of modes $kL/2\pi\in[8.5,25]$. Column (2): The velocity power spectrum. Column (3): The incompressible velocity mode power spectrum. Column (4): The compressible velocity mode power spectrum. Column (5): The kinetic energy power spectrum. Column (6): The sound speed power spectrum.
In Table 3 we show the power law indexes derived by fitting to the scaling ranges of each of the spectra discussed in Section 4 and plotted in Figure 7. For the fits, we use least-squares in $\log-\log$, propagating the $1\sigma$ from the temporal averaging for each of the spectra. As shown in Beattie et al. (2025c), it is not obvious that there is a true inertial (constant $u^{3}/\ell$) range of modes, so we therefore just pick the range $kL/2\pi\in[8.5,25]$ by eye, and perform the fit over that range. Because these simulations are not scale-free, i.e., the box is stratified with scale heights in all quantities (as we show in Figure 3 and Figure 9), we are unable to directly compare the indexes between the galaxy models because $kL/2\pi\in[8.5,25]$ is probing different physical scales in each of the simulations (see Figure 8 for how to properly combine and compare the spectra across galaxies). However, for the same galactic models, with different seeds, we see broad agreement within $1\sigma$, as expected qualitatively from Figure 7.
Appendix C Divergence and Compressible Mode Power Spectrum
In Section 4 we show the compressible mode $\bm{u}_{c}$ spectrum was directly related to the sound speed spectrum, which was controlling the ISM phase structure in our models with a thick disk (Milky Way analogue models). Here we formalize the connection with $\bm{u}_{c}$ and the fluid divergence. The divergence $\bm{\nabla}\cdot\bm{u}$ and $\bm{u}_{c}$ are intrinsically related because only $\bm{u}_{c}$ satisfies $\bm{\nabla}\cdot\bm{u}\neq 0$. In Figure 10 we plot the power spectrum of both quantities (left MW models, right SB models), with the expectation that $du_{c}^{2}(k)/dk\propto d(\bm{\nabla}\cdot\bm{u})^{2}(k)/(k^{2}dk)$ based on dimensionality. Indeed, we find that $du_{c}^{2}(k)/dk\propto d(\bm{\nabla}\cdot\bm{u})^{2}(k)/(k^{2}dk)$ across all resolved modes in the simulations. As we show in Figure 7, $du_{c}^{2}(k)/dk\propto k^{-2}$, which we associate with Burgers (1948) turbulence (as in Beattie et al. 2025a on large scales).
@@ Appendix C Divergence and Compressible Mode Power Spectrum
Figure 11: One-dimensional, isotropic velocity power spectra, $du^{2}(k)/dk$, normalized by the integrated power, as a function of wavenumber, $k$, normalized by the system scale wavenumber, $2\pi/L$. We compensate the spectra by $k^{-5/3}$ to compare to the Kolmogorov 1941 power spectrum. MW runs are shown on the left panel and SB on the right, with each color indicating a different grid resolution of the FX (pinks) and SC (purples) seeding schemes.
Appendix D Convergence test